Explain the structure and geometry of SnCl2 (draw a lewis structure). Estimate the Cl-Sn-Cl bond angles (in gas phase) and explain any deviations from geometrical ideals. Is SnCl2 more stable or less stable than SiCl2, rationalize your answer.
Answer #1

Lewis Structure, Snele suht zu Stre - 4+2 I = 6 electronto = 6/2= 3 election pair sp² hybridisation. . Spe means triangular PSnela is more stable than sich. Bith sif so are members of carbon family, wow, valancy of si fsn is 4. kut due to inestbair e